Find an equation of the line parallel to the graph of 4x+2y=8 and containing the ordered pair (-1,5)
Answer by mananth(16946)

4 x + 2 y = 8
Find the slope of this line
make y the subject
2 y = -4 x + 8
Divide by 2
y = -2 x + 4
Compare this equation with y=mx+b
slope m = -2
The slope of a line parallel to the above line will be the same
The slope of the required line will be -2
m= -2 ,point (-1,5)
Find b by plugging the values of m & the point in
y=mx+b
5 = 2 + b
b= 3
m= -2
Plug value of the slope and b in y = mx +b
The required equation is y=-2x+ 3
m.ananth@hotmail.ca